A vessel containing an unknown volume of argon gas in cm cube at 300 degree Fahrenheit and 101kpa was inverted and placed in cold ethanol.As the gas contracted, the pressure was quadrupled while the temperature reduced by one-third. What is the final volume of the gas in dm3?
1
Expert's answer
2019-04-29T03:15:07-0400

Solution:

pV=nRT

nR is unchanged; nR=1; T1=422 K; T2=300/3=100 Fahrenheit=310,9 K.

V1=1*422/101000=0,0042 L

V2=1*310,9/404000=0,00077 L

Answer: 0,00077 L